OpenStudy (anonymous):

help! what is the simplified form of (15xy^2/x^2-5x+6) / (5x^2y/2x^2-7x+3) ?

OpenStudy (anonymous):

Al-righty here... \[\huge\frac{\frac{15xy^2}{x^2 - 5x + 6}}{\frac{5x^2y}{2x^2 - 7x + 3}} \rightarrow \frac{15xy^2}{x^2 - 5x + 6} \times \frac{2x^2 - 7x + 3}{5x^2y}\]Can you factor the two trinomials?